# methodsReferences for Methods

### v1

@@ Introduction
Multimodeling
Reasonable models
@@
One common approach for combining forecasts is Bayesian model averaging (BMA) [4, 5, 6]. Here, the resulting prediction is given by a mixture of individual models’ posterior predictive distributions where the BMA model weights reflect the fit of a statistical model to data independently of the set of available models. However, BMA relies on theoretical assumptions which are inappropriate for approximate modeling of complex systems (e.g., one of the candidate models is a “true” model that perfectly describes the physical reality).
@@
where $\sigma_{i}$ represents the scale of the error of the mixture model, $\epsilon_{i}\overset{\mathrm{iid}}{\sim}N(0,1)$, $f_{1}(x_{i}),\dots,f_{m}(x_{i})$ are forecasts for the datum $y(x_{i})$ provided by the $m$ theoretical models considered, and $\bm{\omega}(x_{i})\equiv[\omega_{1}(x_{i}),\dots,\omega_{m}(x_{i})]$ are their respective weights which are often adjusted to fulfill the simplex constraint:
$\omega_{1},...,\omega_{m}\geq 0,\leavevmode\nobreak\ \leavevmode\nobreak\ \sum_{k=1}^{m}\omega_{k}=1.$ (2)
@@
Model similarity and redundancy
@@
Model orthogonalization
Principal Component Analysis (PCA)  [16] and Singular Value Decomposition (SVD) [17] are two related methods that have become essential tools for data compression, signal processing, data visualization, feature selection, and dimensionality reduction across science and engineering [18]. In the past [19], PCA has been specifically applied to model orthogonalization; see also other PCA applications to combining forecasts [20, 21].
@@
by subtracting the average $\phi_{0}(x)$ of all the columns,
@@
The vector $\phi_{0}(\bm{x})$, where $\bm{x}=[x_{i}]$ denotes the list of inputs, represents the average forecast of all models and in principle contains the main features that a reasonable physical model should have. The deviations from this average, contained in the matrix $\bm{X^{c}}$, serve to characterize individual models. The matrix $\bm{X^{c}}$ can be expressed in the singular value decomposition form:
$\bm{X^{c}}_{{n\times m}}=\bm{U}_{{n\times n}}\ \ \bm{S}_{{n\times m}}\ \ \bm{V}^{T}_{{m\times m}}\approx\ \ \hat{\bm{X^{c}}}_{{n\times p}}=\hat{\bm{U}}_{{n\times p}}\ \ \hat{\bm{S}}_{{p\times p}}\ \ \hat{\bm{V}}^{T}_{{p\times m}}.$ (6)
In Eq. (6), the reduced-dimension matrix $\hat{\bm{X^{c}}}$ optimally [23, 18] approximates the original matrix $\bm{X^{c}}$ by keeping only the first $p\leq m$ singular values $s_{j}$ of $\bm{S}$. The total number $p$ of components kept can be chosen in several ways (see for example the partial sum criterion [16]), and in this study we treat it as a hyperparameter that is selected by analyzing the performance of the overall model across a validation set.
@@
Once the truncation is done, the retained $p$ principal components are obtained by the columns of $\hat{\bm{U}}$. We label these components as $\phi_{j}(x)$ (with $j=1,\dots,p$). As illustrated in in Fig. 1, the forecast of any of the original models can be approximated by a linear combination of this smaller set of $p$ orthogonal components $\phi_{j}(\bm{x})$ identified by the SVD algorithm, plus the original average forecast:
$f_{k}(\bm{x})\approx\phi_{0}(\bm{x})+\sum_{j=1}^{p}\nu^{(k)}_{j}\phi_{j}(\bm{x}),\ \ \text{for }k=1,\dots,m.$ (7)
@@
Global model mixing and model combination with principal components
@@
$f^{\dagger}(\bm{x};\bm{b})=\sum_{k=1}^{m}\omega_{k}(\bm{b})f_{k}(\bm{x}),$ (9)
@@
where the $m$ weights $\omega_{k}$ depend on the $p$ latent variables $\bm{b}=(b_{1},\dots,b_{p})$ as degrees of freedom. By construction, the sum of the weights $\omega_{k}$ in Eq. (9) adds up to 1, i.e., it satisfies the second part of the simplex constraint (2). Indeed, since $\phi_{0}$ is the average over all models, it contributes with $m\times\frac{1}{m}=1$ to the total sum of the weights, while every principal component $\phi_{j}$ is itself a linear combination of the columns of the matrix $\bm{X^{c}}$, each of which adds net zero total sum of model weights, see Fig. 1. We note, however, that while the weights $\omega_{k}$ fulfill the second simplex constraint, they do not have to be positive.
@@
Given the available data, the weights $\omega_{k}$ can be jointly estimated [24], with the mixture error scale $\sigma$, within a Bayesian framework:
$p(\bm{b},\sigma|\bm{y})\propto p(\bm{y}|\bm{b},\sigma)p(\bm{b},\sigma).$ (10)
@@
Here, $p(\bm{y}|\bm{b},\sigma)$ is the data likelihood function with the standard Gaussian-noise assumption as in Eq. (1), namely
$p(\bm{y}|\bm{b},\sigma)\propto\sigma^{-n}\exp\left(-\frac{1}{2}\chi^{2}\right),\ \ \text{with }\chi^{2}=\sum_{i=1}^{n}\frac{(f^{\dagger}(x_{i};\bm{b})-y(x_{i}))^{2}}{\sigma^{2}},$ (11)
@@
and $p(\bm{b},\sigma)$ is the joint prior distribution of the mixing weights and the common error scale $\sigma$. A reasonable choice for $p(\bm{b},\sigma)$ is multivariate Gaussian prior distribution for $\bm{b}$, informed by the empirical distribution of the original weights $\nu_{j}^{(k)}$ when reproducing the models in (7), and a Gamma prior distribution for the precision $1/\sigma^{2}$ parametrized by a shape parameter $\nu_{0}$ and a scale parameter $\sigma_{0}$, see Ref.[25]. We create a weakly informed prior by selecting $\nu_{0}=10$ and $\sigma_{0}=2$ MeV for the first case study , and $\nu_{0}=10$ and $\sigma_{0}=0.5$ MeV for the realistic case. These choices do not appreciably impact the obtained posterior distributions. Our procedure allows for an efficient approximation of the posterior distribution (10) using the Gibbs sampler[24].
@@
This prior choice effectively resembles BMC in the context of the combined model (9), because the weights $\omega_{k}(\bm{b})$ are unrestricted in the sense that they can be positive and negative. If one wishes to impose a further constraint on the model weights, such as the simplex constraint (2), the original multivariate Gaussian prior distribution for $\bm{b}$ can be easily modified so that it is zero whenever the constraint is not satisfied.
@@
The combined model (8), referred to as BMC+PCA in the following, parameterizes an approximated manifold of the currently developed theoretical models through the mixing weights $\bm{b}$, as illustrated by the grey affine subspace in Fig. 1. In the Bayesian context, this combined model, equipped with a posterior probability distribution for the weights (10), has the appealing statistical interpretation of representing an overarching distribution of plausible models that can explain the observed data. Within this framework, each original forecast $f_{k}$ could be interpreted as a random – not necessarily independent – draw from this distribution, and through Eq. (10) we aim to create a more informed quantified prediction for new observables. Furthermore, performing the mixing on the principal components instead of the original forecasts has the advantage of both filtering out parametric directions that could be associated with noise instead of important features, and prevents the combined-model parameters $\bm{b}$ of becoming ill-conditioned in the presence of similar or redundant forecasts $f_{k}$. We demonstrate these features in the next section.
Case study I: Redundant and Similar Models
@@
${\cal E}(N,Z;\bm{a})=a_{\text{vol}}A+a_{\text{surf}}A^{2/3}+a_{\text{curv}}A^{1/3}+a_{\text{sym}}I^{2}A+a_{\text{ssym}}I^{2}A^{2/3}+a_{\text{sym}}^{(2)}I^{4}A+a_{\rm Coul}{Z^{2}}A^{-1/3},$ (12)
where $A=N+Z$ is the mass number and $I=(N-Z)/A$ is the isospin excess. The parameters $\bm{a}=$ [$a_{\text{vol}}$, $a_{\text{surf}}$, $a_{\text{curv}}$, $a_{\text{sym}}$, $a_{\text{ssym}}$, $a_{\text{sym}}$, $a_{\rm Coul}$] have a well-defined physical meaning; they represent volume, surface, curvature, symmetry, surface-symmetry, second-order-symmetry, and Coulomb terms respectively.
@@
where the model-class index P, G, I, and B, stands for “Perfect", “Good", "Intermediate", and “Bad" models, respectively. These labels reflect how close these models are to the the reference model that generates the synthetic data. Each model class is centered around parameters $\bm{a}^{(t)}$ defined in Table 1. For some scenarios, to obtain non-degenerate forecasts, the parameters are shifted by a small random amount $\delta\bm{a}^{(t)}$, which is a Gaussian with a width of 2‰ of $\bm{a}^{(t)}$. In some cases, we also add a Gaussian noise term with the width $\sigma_{\rm noise}=1$ MeV. These two sources of error: the shift in the parameters and the overall Gaussian noise, simulate a situation in which models within the same class yield predictions that deviate both in a coherent ($\delta\bm{a}^{(t)}$) and uncorrelated ($\sigma_{\rm noise}$) way. (The spread of model predictions due to these sources of noise is shown in Fig. 2(a).) The reference forecast $y_{\rm true}(N,Z)$ consists of $n=629$ binding energies of even-even nuclei with $8\leq Z\leq 102$ computed with the SkO parametrization with the noise term with $\sigma_{\rm noise}$ added.
@@
Table 1: Four model classes used in this study and their respective parameters as defined in Table I of Ref. [28]. The models belonging to the class “Bad" are based on the NL1 parametrization with the terms $\{a_{\text{sym}},a_{\text{ssym}},a_{\text{sym}}\}$ set to zero; we label this parametrization as NL1. Three scenarios (S1-S3) considered are shown on columns 3-5 that list the number $N_{\rm rep,k}$ of repeated (redundant) models belonging to class $k$; the number of principal components $p$ kept; and the individual model weights $\omega_{k}$ (9) obtained by maximizing the likelihood (11) in the case without noise (both $\delta\bm{a}^{(t)}$ and $\sigma_{0}$ are set to zero). Since no noise was added, every repeated model within the same class has identical weight. For each scenario, $\sum\omega_{k}N_{\rm rep,k}=1$. The weights from S1 are correctly re-distributed among the repetitions in S2, i.e., $(\omega_{k}N_{\rm rep,k})_{S1}=(\omega_{k}N_{\rm rep,k})_{S2}$ for each $k$. The perfect model is selected in S3.
We study three scenarios S1-S3, which are described in Table 1 and illustrated in Fig. 2 for scenario S3. In these scenarios, we use different numbers of models in each class with the objective of demonstrating that the proposed algorithm works as intended. The singular values of the SVD (6) can give an initial estimate of the expected number of effective components, as shown Fig. 2(b). The projections of each model on the identified principal components $\phi_{k}$ can visually help identify model classes, as is also shown in the inset of Fig. 2(a).
@@
For the remaining part of this section, we use the BMC+PCA model (8). The synthetic data are separated into three groups: training dataset ($\mathcal{X}_{0}^{tr}$) with 300 datapoints, validation dataset ($\mathcal{X}_{0}^{va}$) with 71 datapoints, and testing dataset ($\mathcal{X}_{0}^{te}$) with 258 datapoints, as is often done in machine learning applications [18], including studies focusing on nuclear mass models [29, 30, 10, 31]. To select the number of principal components kept $p$, we study how it impacts the root mean squared error (RMSE):
$\text{RMSE}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}\Big{(}f^{\dagger}(x_{i};\bm{b}_{0})-y_{\rm true}(x_{i})\Big{)}^{2}}.$ (14)
@@
The parameters $\bm{b}_{0}$ are chosen to maximize the likelihood function (11) on the training dataset $\mathcal{X}_{0}^{tr}$. The RMSE is then computed across all the sets and the number of components $p$ is chosen based on the performance on the validation set. The inset in Fig. 2(b) shows RMSE($p$) for S3. The RMSE computed for the three datasets saturate after $p=3$, as expected. Indeed, there are only four distinct model classes in S3. The training RMSE is always lower, since the parameters $\bm{b}_{0}$ are fitted to it, and the validation RMSE correctly serves as a proxy for the expected RMSE in the extrapolated testing dataset.
Case study II: Realistic nuclear mass models

### v2

@@ I Introduction
II Multimodeling
II.1 Reasonable models
@@ II.2 Combining forecasts
One common approach for combining forecasts is Bayesian model averaging (BMA)  [4, 5, 6]. Here, the resulting prediction is given by a mixture of individual models’ posterior predictive distributions where the BMA model weights reflect the fit of a statistical model to data independently of the set of available models and are obtained by marginalizing over model parameters (i.e., using Bayesian evidence). However, BMA relies on theoretical assumptions which are inappropriate for approximate modeling of complex systems (e.g., one of the candidate models is a “true” model that perfectly describes the physical reality).
@@
where $\sigma_{i}$ represents the scale of the error of the mixture model (possibly including experimental and theoretical errors), $\epsilon_{i}\overset{\mathrm{iid}}{\sim}N(0,1)$, $f_{1}(x_{i}),\dots,f_{m}(x_{i})$ are forecasts for the datum $y(x_{i})$ provided by the $m$ theoretical models considered, and $\bm{\omega}(x_{i})\equiv[\omega_{1}(x_{i}),\dots,\omega_{m}(x_{i})]$ are their respective weights which are often adjusted to fulfill the simplex constraint:
$\omega_{1},...,\omega_{m}\geq 0,~{}~{}\sum_{k=1}^{m}\omega_{k}=1.$ (2)
@@
II.3 Model similarity and redundancy
@@
III Model orthogonalization
Principal Component Analysis (PCA)  [16] and Singular Value Decomposition (SVD) [17] are two related methods that have become essential tools for data compression, signal processing, data visualization, feature selection, and dimensionality reduction across science and engineering [18]. In the past [19], PCA has been specifically applied to model orthogonalization; see also other PCA applications to combining forecasts [20, 21], including a recent application to nuclear mass models [22].
@@ III Model orthogonalization
by subtracting the average $\phi_{0}(x)$ of all the models (columns),
@@
The vector $\phi_{0}(\boldsymbol{x})$, where $\boldsymbol{x}=[x_{i}]$ denotes the list of inputs, represents the average forecast of all models and in principle contains the main features that a reasonable physical model should have. The deviations from this average, contained in the matrix $\bm{X^{c}}$, serve to characterize individual models. The matrix $\bm{X^{c}}$ can be expressed in the singular value decomposition form:
$\bm{X^{c}}_{{n\times m}}$ $=\bm{U}_{{n\times n}}\ \ \bm{S}_{{n\times m}}\ \ \bm{V}^{T}_{{m\times m}}$
$\approx\ \ \hat{\bm{X^{c}}}_{{n\times p}}=\hat{\bm{U}}_{{n\times p}}\ \ \hat{\bm{S}}_{{p\times p}}\ \ \hat{\bm{V}}^{T}_{{p\times m}}.$ (6)
In Eq. (III), the reduced-dimension matrix $\hat{\bm{X^{c}}}$ optimally [24, 18] approximates the original matrix $\bm{X^{c}}$ by keeping only the first $p\leq m$ singular values $s_{j}$ of $\bm{S}$. The total number $p$ of components kept can be chosen in several ways (see for example the partial sum criterion [16]), and in this study we treat it as a hyperparameter that is selected by analyzing the performance of the overall model across a validation set.
@@
Once the truncation is done, the retained $p$ principal components are obtained by the columns of $\hat{\bm{U}}$. We label these components as $\phi_{j}(x)$ (with $j=1,\dots,p$). As illustrated in in Fig. 1, the forecast of any of the original models can be approximated by a linear combination of this smaller set of $p$ orthogonal components $\phi_{j}(\boldsymbol{x})$ identified by the SVD algorithm, plus the original average forecast:
$f_{k}(\boldsymbol{x})\approx\phi_{0}(\boldsymbol{x})+\sum_{j=1}^{p}\nu^{(k)}_{j}\phi_{j}(\boldsymbol{x}),\ \ \text{for }k=1,\dots,m.$ (7)
@@
III.1 Global model mixing and model combination with principal components
@@
$f^{\dagger}(\bm{x};\boldsymbol{b})=\sum_{k=1}^{m}\omega_{k}(\boldsymbol{b})f_{k}(\bm{x}),$ (9)
where the $m$ weights $\omega_{k}$ depend on the $p$ latent variables $\boldsymbol{b}=(b_{1},\dots,b_{p})$ as degrees of freedom:
$\boldsymbol{\omega}(\boldsymbol{b})=\frac{1}{m}\mathbf{1}+\Big{(}\hat{\bm{V}}_{{m\times p}}\hat{\bm{S}}^{-1}_{{p\times p}}\Big{)}\boldsymbol{b},$ (10)
where $\mathbf{1}$ is the all-ones vector of dimension $m$ with all elements equal to 1.
@@
By construction, the sum of the weights $\omega_{k}$ in Eq. (9) adds up to 1, i.e., it satisfies the second part of the simplex constraint (2). Indeed, since $\phi_{0}$ is the average over all models, it contributes with $m\times\frac{1}{m}=1$ to the total sum of the weights, while every principal component $\phi_{j}$ is itself a linear combination of the columns of the matrix $\bm{X^{c}}$, each of which adds net zero total sum of model weights, see Fig. 1. We note, however, that while the weights $\omega_{k}$ fulfill the second simplex constraint, they do not have to be positive.
@@
Given the available data, the weights $\boldsymbol{b}$ can be jointly estimated  [25], with an assumed combined model error scale $\sigma$, within a Bayesian framework:
$p(\boldsymbol{b},\sigma|\boldsymbol{y})\propto p(\boldsymbol{y}|\boldsymbol{b},\sigma)p(\boldsymbol{b},\sigma).$ (11)
@@
Here, $p(\boldsymbol{y}|\boldsymbol{b},\sigma)$ is the data likelihood function with the standard Gaussian-noise assumption as in Eq. (1), namely
$p(\boldsymbol{y}|\boldsymbol{b},\sigma)$ $\propto\sigma^{-n}\exp\left(-\frac{1}{2}\chi^{2}\right),$
$\chi^{2}$ $=\sum_{i=1}^{n}\frac{(f^{\dagger}(x_{i};\boldsymbol{b})-y(x_{i}))^{2}}{\sigma^{2}},$ (12)
and $p(\boldsymbol{b},\sigma)$ is the joint prior distribution of the mixing weights and the common error scale $\sigma$.
The assumption that the deviations between our model predictions $f^{\dagger}(x_{i};\boldsymbol{b})$ and the observations $y(x_{i})$ follow independent Gaussian distributions with the same noise scale ($\sigma_{i}\equiv\sigma$) is a common choice for calibrating nuclear models with extremely precise data such as nuclear masses [26, 27, 28]. This assumption can be easily modified for different applications of the framework if needed.
@@
A reasonable choice for $p(\boldsymbol{b},\sigma)$ is multivariate Gaussian prior distribution for $\boldsymbol{b}$, informed by the empirical distribution of the original weights $\nu_{j}^{(k)}$ when reproducing the models in (7), and a Gamma prior distribution for the precision $1/\sigma^{2}$ parametrized by a shape parameter $\nu_{0}$ and a scale parameter $\sigma_{0}$, see Ref. [29]. We create a weakly informed prior by selecting $\nu_{0}=10$ and $\sigma_{0}=2$ MeV for the first case study, and $\nu_{0}=10$ and $\sigma_{0}=0.5$ MeV for the realistic case. These choices do not appreciably impact the obtained posterior distributions. Notably, our procedure allows for an efficient approximation of the posterior distribution (11) using the standard Gibbs sampler, because the multivariate Gaussian and Gamma priors $p(\boldsymbol{b},\sigma)$ together with the likelihood (III.1) form the classical normal-inverse-gamma semiconjugate model (see Chapter 9 of [29]).
@@
This prior choice effectively resembles BMC in the context of the combined model (9), because the weights $\omega_{k}(\boldsymbol{b})$ are unrestricted in the sense that they can be positive and negative. If one wishes to impose a further constraint on the model weights, such as the simplex constraint (2), the original multivariate Gaussian prior distribution for $\boldsymbol{b}$ can be easily modified so that it is zero whenever the constraint is not satisfied.
@@
The combined model (8), referred to as BMC+PCA in the following, parameterizes an approximated manifold of the currently developed theoretical models through the mixing weights $\boldsymbol{b}$, as illustrated by the grey affine subspace in Fig. 1. In the Bayesian context, this combined model, equipped with a posterior probability distribution for the weights (11), has the appealing statistical interpretation of representing an overarching distribution of plausible models that can explain the observed data. Within this framework, each original forecast $f_{k}$ could be interpreted as a random – not necessarily independent – draw from this distribution, and through Eq. (11) we aim to create a more informed quantified prediction for new observables. Furthermore, performing the mixing on the principal components instead of the original forecasts has the advantage of both filtering out parametric directions that could be associated with noise instead of important features, and prevents the combined-model parameters $\boldsymbol{b}$ of becoming ill-conditioned in the presence of similar or redundant forecasts $f_{k}$. We demonstrate these features in the next section.
IV Results
IV.1 Case Study I: Redundant and similar models
@@
${\cal E}(N,Z;\boldsymbol{a})$ $=a_{\text{vol}}A+a_{\text{surf}}A^{2/3}+a_{\text{curv}}A^{1/3}+a_{\text{sym}}I^{2}A$
$\quad+a_{\text{ssym}}I^{2}A^{2/3}+a_{\text{sym}}^{(2)}I^{4}A+a_{\rm Coul}\frac{Z^{2}}{A^{1/3}},$ (13)
where $A=N+Z$ is the mass number and $I=(N-Z)/A$ is the isospin excess. The parameters $\boldsymbol{a}=$ [$a_{\text{vol}}$, $a_{\text{surf}}$, $a_{\text{curv}}$, $a_{\text{sym}}$, $a_{\text{ssym}}$, $a_{\text{sym}}$, $a_{\rm Coul}$] have a well-defined physical meaning; they represent volume, surface, curvature, symmetry, surface-symmetry, second-order-symmetry, and Coulomb terms respectively.
@@
where the model-class index P, G, I, and B, stands for “Perfect", “Good", "Intermediate", and “Bad" models, respectively. These labels reflect how close these models are to the the reference model that generates the synthetic data. Each model class has parameters centered around parameters $\boldsymbol{a}^{(t)}$ defined in Table 1. For some scenarios, to obtain non-degenerate forecasts, the parameters are shifted by a small random amount $\delta\boldsymbol{a}^{(t)}$, which is a Gaussian with a width of 2‰ of $\boldsymbol{a}^{(t)}$. In some cases, we also add a Gaussian noise term with the width $\sigma_{\rm noise}=1$ MeV. These two sources of error: the shift in the parameters and the overall Gaussian noise, simulate a situation in which models within the same class yield predictions that deviate both in a coherent ($\delta\boldsymbol{a}^{(t)}$) and uncorrelated ($\sigma_{\rm noise}$) way. (The spread of model predictions due to these sources of noise is shown in Fig. 2(a).) The reference forecast $y_{\rm true}(N,Z)$ consists of $n=629$ binding energies of even-even nuclei with $8\leq Z\leq 102$ computed with the SkO parametrization with the noise term with $\sigma_{\rm noise}$ added.
Table 1: Four model classes used in this study and their respective parameters as defined in Table I of Ref.  [32]. The models belonging to the class “Bad" are based on the NL1 parametrization with the terms $\{a_{\text{sym}},a_{\text{ssym}},a^{(2)}_{\text{sym}}\}$ set to zero; we label this parametrization as NL1. Three scenarios (S1-S3) considered are shown on columns 3-5 that list the number $N_{\rm rep,k}$ of repeated (redundant) models belonging to class $k$; the number of principal components $p$ kept; and the individual model weights $\omega_{k}$ (9) obtained by maximizing the likelihood (III.1) in the case without noise (both $\delta\boldsymbol{a}^{(t)}$ and $\sigma_{0}$ are set to zero). Since no noise was added, every repeated model within the same class has identical weight. For each scenario, $\sum\omega_{k}N_{\rm rep,k}=1$. The weights from S1 are correctly re-distributed among the repetitions in S2, i.e., $(\omega_{k}N_{\rm rep,k})_{S1}=(\omega_{k}N_{\rm rep,k})_{S2}$ for each $k$. The perfect model is selected in S3.
@@
We study three scenarios S1-S3, which are described in Table 1 and illustrated in Fig. 2 for scenario S3. In these scenarios, we use different numbers of models in each class with the objective of demonstrating that the proposed algorithm works as intended. The singular values of the SVD (III) can give an initial estimate of the expected number of effective components, as shown Fig. 2(b). The projections of each model on the identified principal components $\phi_{k}$ can visually help identify model classes, as is also shown in the inset of Fig. 2(a).
@@
For the remaining part of this section, we use the BMC+PCA model (8). The synthetic data are separated into three groups: training dataset ($\mathcal{X}_{0}^{tr}$) with 300 datapoints, validation dataset ($\mathcal{X}_{0}^{va}$) with 71 datapoints, and testing dataset ($\mathcal{X}_{0}^{te}$) with 258 datapoints, as is often done in machine learning applications [18], including studies focusing on nuclear mass models [33, 34, 10, 35]. The specific way the three sets for our work was chosen (see Panel (a) of Figure 3) reflects that one of the main objectives of our model forecast combination lies in model extrapolation into the region in which experimental information does not exist. By dividing the sets in this way, we provide a more stringent test of how the combined model’s performance will evolve as we go further away from the region where data currently exists.
To select the number of principal components kept $p$, we study how it impacts the root mean squared error (RMSE):
$\text{RMSE}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}\Big{(}f^{\dagger}(x_{i};\boldsymbol{b}_{0})-y_{\rm true}(x_{i})\Big{)}^{2}}.$ (15)
@@
The parameters $\boldsymbol{b}_{0}$ are chosen to maximize the likelihood function (III.1) on the training dataset $\mathcal{X}_{0}^{tr}$. The RMSE is then computed across all the sets and the number of components $p$ is chosen based on the performance on the validation set. The inset in Fig. 2(b) shows RMSE($p$) for S3. The RMSE computed for the three datasets saturate after $p=3$, as expected. Indeed, there are only four distinct model classes in S3. The training RMSE is always lower, since the parameters $\boldsymbol{b}_{0}$ are fitted to it, and the validation RMSE correctly serves as a proxy for the expected RMSE in the extrapolated testing dataset.
IV.2 Case Study II: Realistic nuclear mass models